Thin films mornng report 4/30/04

james haydon jhaydon at snf.stanford.edu
Fri Apr 30 09:37:00 PDT 2004


           

Thinfilms Morning Report 4/30/2004

 

 

Gryphon:            ok

Innotec;            ok

Balzers :            ok

STS Dep:            ok

Implanter:            phosphine bottle change

Metalica:            cryo pump replacement 5/3/04

SCT:                completion of installation near end of May

CPD:                ok                   

WbGen-CTB:            control temp bath not working should be back up 
late next week

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            right hot pot continues to drain after drain has 
been turned off

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok

 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/thinfilms/attachments/20040430/2927cfa0/attachment.html>


More information about the thinfilms mailing list